Project

General

Profile

Kaon LT Meetings » mtg_22mar23.txt

Richard Trotta, 03/24/2022 02:05 PM

 
1

    
2
		Mar 23/22 PionLT/KaonLT Analysis Meeting Notes
3
                ----------------------------------------------
4

    
5
Participants:
6
 - Richard Trotta, CUA
7
 - Pete Markowitz, FIU
8
 - Dave Gaskell, JLab
9
 - Julie Roche, Ohio
10
 - Nathan Heinrich, Garth Huber, Muhammad Junaid, Stephen Kay, Vijay Kumar,
11
   Love Preet, Ali Usman, Regina
12

    
13
Please remember to post your slides at:
14
https://redmine.jlab.org/projects/kltexp/wiki/Kaon_LT_Meetings
15

    
16
Richard
17
-------
18
- 10.6 GeV Heep-COIN analysis and SIMC comparison
19
  - lots of crashes when analyzing Heep-singles, so nothing on that yet
20
  - comparison includes Comp-LT instead of EDTM-LT
21
  - tracking cuts not yet implemented.  Tracking efficiencies are included in
22
    the Heep-COIN comparison, but won't be accurate until cuts are checked
23
  - Data/SIMC discrepancy is about 10%, so more work is needed
24
  - standard.kinematics not yet double-checked (see Ali report)
25

    
26
- *Suggestions*
27
  GH: please show intermediate plots next time, so we can more easily see
28
      what was done (as well as to provide documentation)
29
    - a variety of DAQ and other issues were encountered when KaonLT started up
30
      at 10.6 GeV, such as incorrect CoinTime setup.  Please be sure to take a
31
      close look at these variables, so we can see what improvements need to be
32
      made to the analysis
33

    
34
- *Goal for Next Meeting*
35
  - improved 10.6 GeV Heep-COIN and Heep-Singles plots (data and simulations)
36

    
37
Ali
38
---
39
- using LTSEP_2022 analysis branch
40
  - started replaying, crashing a lot
41
  - param and other files were in the wrong place
42
  - created a specific directory for KaonLT, rearranged files and replay is
43
    working now
44
  - doing some full replays to be sure everything is working
45

    
46
- checking standard.kinematics for KaonLT higher energy runs
47
  - Arc energy measurements not all correct, particularly for 10.6 GeV
48

    
49
- met w/Richard to coordinate, and another meeting planned in a few days
50
  - Richard will continue to work on the analysis structure
51
  - Ali will look at angles and momenta in standard.kinematics for all higher
52
    energy settings
53

    
54
- *Goal for Next Meeting*
55
  - 8.6 GeV Heep-COIN plots (data and simulations) hopefully with efficiencies,
56
    Comp-LT and improved tracking
57

    
58
Vijay
59
-----
60
- CTime_eKcointime plot from 3.8 GeV data
61
  - multiple peaks were shown at last meeting, worked on identifying them
62
  - peak to right is protons, confirmed two ways:
63
    1 calculated the TOF difference for pi/K/protons.  The peak spacing is
64
      consistent with his calculation (in logbook only, please add to slides
65
      before posting them)
66
    2 looked at MM versus epcointime.  proton band is straight, while it was
67
      tilted in eKcointime vs MM
68
  - this means the line (in MM vs cointime) near the neutron MM is actually
69
    omega peak
70

    
71
- Data/MC comparison (SIMC x7) at 3.8 GeV
72
  - still no efficiencies included
73
  - saw some tails in the SIMC SHMS distributions not evident in the data.
74
  - In the discussion, it was realized there was an MMK cut to select Lambda
75
    peak that was applied to the data, but not SIMC.
76
  - It is important to have the same cuts in data and SIMC, so applying this
77
    cut will probably get rid of the tails in the SIMC distributions.  Vijay
78
    will add this.
79

    
80
- checking standard.kinematics for KaonLT and PionLT low energy runs
81
  - Arc energy measurements mostly correct, although some round-off errors need
82
    to be fixed
83
  - spectrometer angles/momenta not yet checked
84

    
85
- *Goal for Next meeting*
86
  - 4.9 GeV data/SIMC comparison for physics settings
87

    
88
- *Suggestions*
89
  DG: It is extremely important to do the standard.kinematics double-check
90
      BEFORE doing the data/MC comparison, as otherwise you're wasting time
91
    - Brad has a script to extract the spectrometer photos from CODA.  It would
92
      be good to get this script from him, as KaonLT record keeping was
93
      sometimes inconsistent, and it would save time having to check through a
94
      lot of hclogs for the information
95

    
96
Nathan
97
------
98
- Mode-10 Data Analysis Update
99
  - at last meeting, there was a discrepancy betwen hardware and software
100
    versions of the Pulse Amplitude
101
  - Mark suggested this could be due to phger_SampNSAT variable, which is 1mV
102
    for hardware but 2mV default for software
103
  - Nathan set the threshold to 1mV, now the agreement is very good
104
  - get ~1% more events with 1mV threshold (10k out of 1.2M evts)
105
  - Mark's implementation looks good for hgcer
106
    - has now added the Aerogel for Mode10 runs
107

    
108
- *Goal for Next Meeting*
109
  - Look at the NGC and Aerogel plots to see how they look compared to data
110
  - Then wants to look into the HGC Zero Amplitude events issue that Vijay
111
    sees.  Plan is to select HGC Zero Amplitude events and see what the
112
    software variables give for these events
113

    
114
- *Reminder of task for DG*
115
  - DG will talk to Mark/Brad on whether we can get the FADC firmware fixed to
116
    return BOTH firmware and waveform data when there is an error in time for
117
    the upcoming run
118

    
119
Stephen
120
-------
121
- was helping Vijay and Ali with various replay issues
122
  - hcana update: now pointing to Mark's firmware branch to include the Mode-10
123
    data.  This is not the default branch
124
  - no major ROOT version issues encountered yet
125
    - but did need to turn off CUE (Common Unix Environment) initialization, as
126
      this causes a c++ compiler incompatibility version between hcana and ROOT
127

    
128
- batch queuing
129
  - using swif, but finds it harder to see what's going on with job
130
    progress/crashes compared to auger
131

    
132
- software for new run
133
  - would like to add the Mode-10 variables and create a new hcana version for
134
    online
135
  - python environment should be okay
136

    
137
Jacob (not present)
138
-----
139
- *Task for Jacob*
140
  - Jacob will follow up with DM and CY about the status of the calibration
141
    code.  It would be great if we could test it on the PionLT and KaonLT data,
142
    and see if it resolves any of the hodoscope calibration issues we have
143

    
144

    
145
Upcoming Run
146
------------
147
- GH: we need to start the planning of the Heep and Lumi studies at our next
148
  meeting
149
- DG: has been attending the 08:00 meetings.
150
  - Things still on track for the June run as scheduled
151
  - Linac gradient will not be known for some time yet, but nothing has broken
152
    yet, so we should continue planning with the expected energies
153
  - will start with the same target configuration as the fall run
154

    
155
Next Meeting
156
------------
157
- Wednesday April 6 at 13:00 Eastern/11:00 Regina
158
- after that, GH will circulate a poll to select a new time (maybe Tue or Thu)
159
  to avoid Tanja's teaching conflict
(35-35/413)